# Transverse response from anisotropic Fermi surfaces

### v2

@@ I Introduction
In general, for an isotropic 2DEG with a circular Fermi surface, transverse currents cancel because states with opposite transverse momenta carry equal-and-opposite contributions. However, this cancellation breaks down when the crystallographic axes of a 2DEG with an anisotropic Fermi surface are rotated relative to the transport direction. In such a configuration, states with opposite transverse momenta and identical longitudinal momentum need not exist, and a net transverse current can appear even without magnetic fields or broken time-reversal symmetry. Such an effect rooted in anisotropy of the Fermi contour has been studied in the context of Josephson junctions [9, 10].
In this work, we analyse this effect in detail. We first compute the transverse conductivity of a translationally invariant system with an anisotropic Fermi surface. We then examine the emergence of a transverse voltage in a lattice model of a 2DEG using the Büttiker-probe technique [11]. It is known that Hall voltage can appear in absence of time-reversal symmetry breaking due to reflection [12] or refraction [13] in the sample due to potential profile. Our results demonstrate that anisotropy and crystalline orientation is yet another way to generate transverse responses conventionally associated with Hall phenomena, despite the complete absence of magnetic fields or any mechanism that breaks time-reversal symmetry.
@@ II Continuum model
$H_{c}$ $=$ $-(t-\delta\cos{2\phi})a^{2}\partial_{x}^{2}-(t+\delta\cos{2\phi})a^{2}\partial_{y}^{2}$ (1)
$+2\delta\sin{2\phi}~a^{2}\partial_{x}\partial_{y},$
where $\delta$ denotes the anisotropy parameter and $\phi$ specifies the angle between the major axis of the elliptical Fermi contour and the $\hat{x}$ direction, and $a$ is a length-scale that can be thought of as lattice spacing of the underlying lattice. For simplicity, we choose $0\leq\delta<t$. In the limiting case $\delta=0$, the Fermi contour becomes isotropic, corresponding to an effective mass $\hbar^{2}/2t$.
@@
$J_{x}$ $=$ $2e[(t-\delta\cos{2\phi})k_{x}-\delta\sin{2\phi}~k_{y}]a^{2}/\hbar,$
$J_{y}$ $=$ $2e[(t+\delta\cos{2\phi})k_{y}-\delta\sin{2\phi}~k_{x}]a^{2}/\hbar.$ (2)
@@ III Lattice Model
Figure 4: (a) Transverse voltage as a function of $\phi$, the angle of rotation of the system. Parameters: $L_{x}=L_{y}=11$, $n_{x0}=6$, $\mu_{n}=-2t$, $\mu_{M}=-4t$, $t^{\prime}=t$, $t_{P}=0.05t$, and $eV=0.05t$. Different curves correspond to different values of $\delta$. (b) Fermi contour of the central lattice for various choices of $\delta$ at energy $0.1t$. The legends indicate the values of $\delta/t$.
We now briefly outline the method of calculation. The scattering eigenfunction corresponding to an electron incident in the $m_{0}$-th channel of the source at energy $E$ is first obtained using standard scattering theory. This eigenfunction is then used to compute the currents flowing into the probe terminals. The total current through each probe terminal due to a longitudinal bias $V$ is obtained by summing the contributions from all incident channels within the bias window. Next, scattering eigenfunctions for electrons incident from each of the probe terminals are computed, and the corresponding currents are evaluated for a wide range of voltage pairs $(V_{T},V_{B})$ applied at the probe terminals. The total currents in the probe terminals thus include contributions from the source–drain bias $V$ as well as from the applied probe-terminal voltages. From the set of sampled voltage pairs, we identify those for which the net current in both probe terminals vanishes. The corresponding voltage difference $V_{T}-V_{B}$ defines the transverse voltage $V_{H}$. All calculations are performed numerically.
@@
IV Discussion and Conclusion

### v3

@@ I Introduction
In general, for an isotropic 2DEG with a circular Fermi surface, transverse currents cancel because states with opposite transverse momenta carry equal and opposite contributions. However, this cancellation breaks down when the crystallographic axes of a 2DEG with an anisotropic Fermi surface are rotated relative to the transport direction. In such a configuration, states with opposite transverse momenta and identical longitudinal momentum need not exist, and a net transverse current can arise even in the absence of magnetic fields or broken time-reversal symmetry. Effects rooted in the anisotropy of the Fermi contour have previously been discussed in the context of Josephson junctions [12, 11].
In this work, we analyse this effect in detail. We first compute the transverse conductivity of a translationally invariant system with an anisotropic Fermi surface. We then examine the emergence of a transverse voltage in a lattice model of a 2DEG using the Büttiker-probe technique [21]. While the continuum model yields a transverse conductivity in a translationally invariant setting, the lattice model probes the corresponding transverse voltage in a multiterminal geometry; the latter provides an experimentally relevant manifestation of the same underlying anisotropic transport mechanism. It is known that transverse voltages can arise even in the absence of time-reversal symmetry breaking due to reflection [18] or refraction [14] induced by spatial inhomogeneities in the potential landscape. Our results demonstrate that anisotropy and crystalline orientation provide a distinct route to generating transverse responses, without relying on magnetic fields, Berry curvature, or spatially varying potentials.
It is useful to contrast the present mechanism with related effects discussed in the literature. In systems with Rashba spin-orbit coupling subject to an in-plane Zeeman field, the Fermi contour is displaced away from $(k_{x},k_{y})=(0,0)$, which leads to an imbalance between states with opposite transverse momenta; this underlies the planar Hall effect and related transverse responses in Josephson systems [20, 12]. In contrast, in the work of Song et al. [18], the transverse response arises from asymmetric reflection of electrons by engineered antidot structures, while in the work of Sarkar et al. [14], it originates from refraction-like bending of electron trajectories due to spatial variations in the potential. In all these cases, the transverse response is tied either to a shift of the Fermi surface or to real-space inhomogeneities. In contrast, the mechanism discussed here relies solely on a rotated anisotropic Fermi surface that remains centred at $(k_{x},k_{y})=(0,0)$, and therefore does not require any symmetry breaking in real space or momentum-space displacement of the Fermi contour.
@@ II Continuum model
$H_{c}$ $=$ $-(t-\delta\cos{2\phi})a^{2}\frac{\partial^{2}}{\partial x^{2}}-(t+\delta\cos{2\phi})a^{2}\frac{\partial^{2}}{\partial y^{2}}$ (1)
$+2\delta\sin{2\phi}~a^{2}\frac{\partial^{2}}{\partial x\partial y},$
where $\delta$ denotes the anisotropy parameter and $\phi$ specifies the angle between the major axis of the elliptical Fermi contour and the $\hat{x}$ direction, and $a$ is a length-scale that can be thought of as lattice spacing of the underlying lattice. Here, $\partial/\partial x$ denotes the spatial derivative, corresponding to the momentum operator $i\hbar\partial_{x}$ in the continuum representation. For simplicity, we choose $0\leq\delta<t$. In the limiting case $\delta=0$, the Fermi contour becomes isotropic, corresponding to an effective mass $\hbar^{2}/2ta^{2}$.
@@
$J_{x}$ $=$ $2e[(t-\delta\cos{2\phi})k_{x}-\delta\sin{2\phi}~k_{y}]/\hbar,$
$J_{y}$ $=$ $2e[(t+\delta\cos{2\phi})k_{y}-\delta\sin{2\phi}~k_{x}]/\hbar.$ (2)
@@ III Lattice Model
Physically, the parameter $\delta$ quantifies the anisotropy between the effective hopping amplitudes along two orthogonal directions, while $\phi$ denotes the angle between the principal axes of the anisotropic bandstructure and the transport direction. To estimate realistic values of $\delta$, we relate it to experimentally measured transport anisotropy. From Eq. (3), the longitudinal conductivities at $\phi=0$ and $\phi=\pi/2$ satisfy
$\frac{G_{xx}(\phi=0)}{G_{xx}(\phi=\pi/2)}=\sqrt{\frac{t-\delta}{t+\delta}}.$
Using reported anisotropic resistance data for CrSBr, a material known to host a strongly anisotropic bandstructure [16], we obtain an estimate $\delta\approx 0.43t$, indicating that the parameter regime considered in this work is experimentally realistic.
More broadly, anisotropic hopping parameters of this kind can arise in a variety of systems, including layered van der Waals materials with low crystal symmetry, and strain-engineered materials [6, 4] where hopping amplitudes and lattice orientation can be tuned in a controlled manner.
@@
Figure 4: Schematic illustrating the Büttiker probe technique. The sample is connected to source and drain terminals along the longitudinal direction, and to top and bottom probe terminals in the transverse direction. A bias applied between source and drain drives a longitudinal current, which is deflected asymmetrically toward the transverse probes due to anisotropy. The probe voltages $V_{T}$ and $V_{B}$ adjust self-consistently such that the net current through each probe vanishes. The resulting voltage difference $V_{H}=V_{T}-V_{B}$ defines the transverse voltage.
We now briefly outline the method of calculation and provide an intuitive picture of the transverse response within the Büttiker probe framework. In a multiterminal setup, the probe terminals are defined such that the net current flowing into them is zero. The probe voltages adjust self-consistently to ensure zero net current in the transverse terminals, thereby preventing charge accumulation. When the system exhibits an imbalance in the flow of carriers toward the two transverse directions, the probe voltages $V_{T}$ and $V_{B}$ shift relative to each other to compensate for this imbalance, resulting in a finite transverse voltage.
We compute this quantitatively using scattering theory. The scattering eigenfunction corresponding to an electron incident in the $m_{0}$-th channel of the source at energy $E$ is first obtained. This eigenfunction is used to evaluate the currents flowing into the probe terminals. The total current through each probe terminal due to a longitudinal bias $V$ is obtained by summing over all incident channels within the bias window.
Next, scattering eigenfunctions corresponding to electrons incident from each of the probe terminals are computed, and the associated currents are evaluated for a range of probe voltages $(V_{T},V_{B})$. The total currents in the probe terminals therefore include contributions from both the source–drain bias $V$ and the applied probe voltages. The values of $(V_{T},V_{B})$ are then determined by imposing the condition that the net current in each probe terminal vanishes. The resulting voltage difference $V_{T}-V_{B}$ defines the transverse voltage $V_{H}$. All calculations are performed numerically.
Figure 5: (a) Transverse voltage as a function of $\phi$, the angle of rotation of the system. Parameters: $L_{x}=L_{y}=11$, $n_{x0}=6$, $\mu_{n}=-2t$, $\mu_{M}=-4t$, $t^{\prime}=t$, $t_{P}=0.05t$, and $eV=0.05t$. Different curves correspond to different values of $\delta$. (b) Fermi contour of the central lattice for various choices of $\delta$ at energy $0.1t$. The legends indicate the values of $\delta/t$. (c) Hall voltage versus system size $L_{x}=L_{y}$ for fixed values of $\phi=\pi/4$, $eV/t=10^{-5}$. The Hall voltage is calculated at the center, i.e., $n_{x0}=(L_{x}+1)/2$.
@@
To examine finite-size effects, in Fig. 5(c) we show the scaling of the transverse voltage $V_{H}$ with system size. We fix the applied bias and compute $V_{H}$ as a function of $L_{x}=L_{y}$, evaluating it at the central site $n_{x0}=(L_{x}+1)/2$. We find that the magnitude of $V_{H}$ does not show systematic suppression with increasing system size, indicating that the effect is not a finite-size effect.
Importantly, the angles $\phi$ at which the transverse response vanishes remain unchanged as the system size is increased. This demonstrates that the symmetry-protected zeros of $V_{H}$ are robust and not affected by finite-size effects.
The dependence of the transverse response on the Fermi energy can be understood analytically within the continuum model. As shown in Eq. (II), the transverse conductivity scales as $\sqrt{E}$, reflecting the energy dependence of the density of states and the underlying anisotropic dispersion. This indicates that the transverse response varies smoothly with carrier density and can, in principle, be tuned continuously by electrostatic gating in experiments. A similar qualitative dependence shows up in the lattice model, since it reproduces the continuum dispersion near the band bottom. Although relatively large biases are used in the numerical calculations for clarity, the transverse voltage scales linearly with the applied bias.
It is instructive to contrast this behavior with that of conventional Hall effects. In the quantum Hall effect, the Hall conductivity exhibits quantized plateaus as a function of Fermi energy due to Landau level formation in the presence of a magnetic field. In contrast, the transverse response in the present work arises purely from Fermi-surface anisotropy and varies continuously with energy, without requiring magnetic fields.
IV Discussion
The emergence of a transverse response in our system can be understood from symmetry considerations. In an isotropic system or when the principal axes of an anisotropic Fermi surface are aligned with the transport direction, the dispersion is symmetric under $k_{y}\to-k_{y}$. As a result, states with opposite transverse momenta contribute equally and oppositely to the transverse current, leading to its cancellation. However, when the anisotropic Fermi surface is rotated with respect to the transport axes, this mirror symmetry is broken, and the cancellation no longer holds, resulting in a finite transverse response.
The transverse conductivity therefore vanishes at special orientations $\phi=n\pi/2$ (with integer $n$), where the mirror symmetry $k_{y}\to-k_{y}$ is restored. We emphasize that this effect does not involve breaking of time-reversal symmetry.
Onsager reciprocity relations apply to the full multiterminal conductance matrix $G_{pq}$, which satisfies $G_{pq}(B)=G_{qp}(-B)$ (where $B$ is magnetic field) and is symmetric at zero magnetic field. Here, the subscripts $p,q$ refer to the indices that represent terminals and $G_{pq}$ refers to the differential ratio of current in terminal $p$ to the voltage in terminal $q$. On the other hand, $G_{yx}$ calculated in our work refers to the differential ratio of current in the $y$-direction to voltage applied in the $x$-direction. The transverse response reported in our work using lattice model is obtained after imposing probe conditions (zero current in the transverse terminals) and therefore corresponds to an effective response coefficient rather than a direct element of the conductance matrix. As a result, it is not constrained to satisfy a simple symmetry under index exchange.
The transverse response discussed here originates from the bulk anisotropy of the dispersion and is therefore expected to be robust against variations in contact properties, although the precise magnitude of the measured voltage may depend on contact details.
Examples of materials exhibiting strongly anisotropic bandstructures include quasi-two-dimensional organic conductors such as $\beta^{\prime\prime}$-(BEDT-TTF) salts [13, 1], van der Waals materials such as CrSBr [16], and low-symmetry transition metal dichalcogenides such as ReSe2 [9, 10]. In particular, CrSBr has recently attracted significant attention due to its pronounced in-plane anisotropy arising from its low-symmetry crystal structure [16], making it a promising platform for realizing the mechanism discussed in this work. These systems provide experimentally accessible platforms where the orientation of anisotropic electronic structure relative to transport direction can be controlled, for example via device geometry or strain.
The results presented here are obtained in the ballistic regime. The underlying mechanism, however, originates from the anisotropy of the Fermi surface and is therefore expected to persist in the presence of weak to moderate disorder. Disorder primarily affects the magnitude of the response, and only in the limit of strong disorder the transverse response can be suppressed.
In the low-filling regime, the lattice model reproduces the continuum dispersion near the band bottom, and therefore the two approaches yield qualitatively consistent results. A direct quantitative comparison between the two approaches is, however, not straightforward, since the continuum calculation yields a transverse conductivity in a translationally invariant system, whereas the lattice calculation determines a transverse voltage in a finite multiterminal geometry using the Büttiker probe framework.
V Conclusion